OpenStudy (anonymous):

How do I solve this? Multiply the complex numbers below (-3+2i) (2+i)

OpenStudy (anonymous):

you know foil? try that

OpenStudy (anonymous):

so (-3 + 2i)(2 + i) -6 - 3i + 4i + 2i^2

OpenStudy (anonymous):

Yes I got that far, but then i get -6-i-2

OpenStudy (anonymous):

you can simplify that a little further

OpenStudy (anonymous):

-6-i-2 = -8 - i

OpenStudy (phi):

you mean -8 + i (not -8 -i)

OpenStudy (anonymous):

oops, that's right! -8 + i. thanks @phi

OpenStudy (anonymous):

-3i+4i=1i i^2=-1

OpenStudy (anonymous):

oh, I was stuck at that part, not sure how to get the answer

OpenStudy (anonymous):

but i got it now :) thank you guys